# Comparing Inferential Strategies of Humans and Large Language Models in Deductive Reasoning

### v1

@@ Appendix C Annotated Model Responses
Within this section, we showcase examples of model responses that exemplify each inferential strategy identified in our study, as depicted in figures 6-19. Each figure is organized with the problem statement at the top, the model’s response on the lower left, and the annotators’ comments to the lower right. For an extensive array of model responses and annotations, we invite readers to explore our data repository, which we publish upon acceptance of this paper.
@@ C.1 Supposition Following
Model |
[…]
green!20 Supposition Following |
[…]
yellow!20 Chain Construction |
[…]
orange!20 Compound Conclusion |
[…]
mylightblue Concatenation Strategy |
[…]
blue!20 Symbolic Strategy |
Correct Answer |
Sound Reasoning |
Zephyr-7B- |
Mistral-7B-Instruct |
LLaMA-2-7B |
LLaMA-2-13B |
LLaMA-2-70B |
@@ C.5 Symbolic Strategy
Model |
Base Model |
Parameters |
Context Length |
Tokens |
Fine-tuning |
Zephyr-7B- |
Mistral |
7B |
8192 tokens |
dSFT, AIF |
Mistral-7B-Instruct |
Mistral |
7B |
8192 tokens |
SFT |
LLaMA-2-7B-Chat |
LLaMA-2 |
7B |
4K tokens |
2.0T |
SFT, RLHF |
LLaMA-2-13B-Chat |
LLaMA-2 |
13B |
4K tokens |
2.0T |
SFT, RLHF |
LLaMA-2-70B-Chat |
LLaMA-2 |
70B |
4K tokens |
2.0T |
SFT, RLHF |

### v2

@@ A.4 Model Details
| Model | Base Model | Parameters | Context Length | Tokens | Fine-tuning |
| Zephyr-7B- | Mistral | 7B | 8192 tokens | - | dSFT, AIF |
| Mistral-7B-Instruct | Mistral | 7B | 8192 tokens | - | SFT |
| LLaMA-2-7B-Chat | LLaMA-2 | 7B | 4K tokens | 2.0T | SFT, RLHF |
| LLaMA-2-13B-Chat | LLaMA-2 | 13B | 4K tokens | 2.0T | SFT, RLHF |
| LLaMA-2-70B-Chat | LLaMA-2 | 70B | 4K tokens | 2.0T | SFT, RLHF |
@@ Appendix C Annotated Model Responses
Within this section, we showcase examples of model responses that exemplify each inferential strategy identified in our study, as depicted in figures 6-19. Each figure is organized with the problem statement at the top, the model’s response on the lower left, and the annotators’ comments to the lower right. For an extensive array of model responses and annotations, we invite readers to explore our data repository at: huggingface.co/datasets/mainlp/inferential_strategies.
| Model | Supposition Following | Chain Construction | Compound Conclusion | Concatenation Strategy | Symbolic Strategy | Correct Answer | Sound Reasoning |
| Zephyr-7B- | |||||||
| Mistral-7B-Instruct | |||||||
| LLaMA-2-7B | |||||||
| LLaMA-2-13B | |||||||
| LLaMA-2-70B |
@@ C.5 Symbolic Strategy
Statements: 1. There is a blue marble in the box if and only if there is a red marble in the box. 2. Either there is a red marble in the box or else there is a white marble in the box, but not both. 3. There is a white marble in the box if and only if there is a pink marble in the box. Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
@@ 1. First, I’ll start by analyzing the given statements:
Statements: 1. There is a blue marble in the box if and only if there is a red marble in the box. 2. Either there is a red marble in the box or else there is a white marble in the box, but not both. 3. There is a white marble in the box if and only if there is a pink marble in the box. Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
@@
Statements: 1. There is a purple marble in the box if and only if there is a blue marble in the box. 2. There is a blue marble in the box if and only if there is an olive marble in the box. 3. Either there is an olive marble in the box or else there is a red marble in the box, but not both. 4. Either there is a red marble in the box or else there is a green marble in the box, but not both. Conclusion: If there is a purple marble in the box then there is a green marble in the box.
@@ 1. First, let’s analyze the given statements one by one:
Statements: 1. There is a purple marble in the box if and only if there is a blue marble in the box. 2. There is a blue marble in the box if and only if there is an olive marble in the box. 3. Either there is an olive marble in the box or else there is a red marble in the box, but not both. 4. Either there is a red marble in the box or else there is a green marble in the box, but not both. Conclusion: If there is a purple marble in the box then there is a green marble in the box.
@@ 1. First, I’ll start by analyzing the given statements:
Statements: 1. There is a blue marble in the box if and only if there is a red marble in the box. 2. Either there is a red marble in the box or else there is a white marble in the box, but not both. 3. There is a white marble in the box if and only if there is a pink marble in the box. Conclusion: If there is not a blue marble in the box then there is a pink marble in the box.
@@
Statements: 1. There is a brown marble in the box if and only if there is a yellow marble in the box. 2. Either there is a yellow marble in the box or else there is a green marble in the box, but not both. 3. There is a green marble in the box if and only if there is a purple marble in the box. 4. There is a purple marble in the box if and only if there is an olive marble in the box. Conclusion: If there is a brown marble in the box then there is an olive marble in the box.
@@ 4. There is a purple marble in the box if and only if there is an olive marble in the box
Statements: 1. There is a purple marble in the box if and only if there is a yellow marble in the box. 2. There is a yellow marble in the box if and only if there is a blue marble in the box. 3. Either there is a blue marble in the box or else there is an orange marble in the box, but not both. Conclusion: Either there is a purple marble in the box or else there is an orange marble in the box, but not both.
@@
Statements: 1. There is a blue marble in the box if and only if there is a brown marble in the box. 2. There is a white marble in the box if and only if there is a green marble in the box. 3. Either there is a brown marble in the box or else there is a white marble in the box, but not both. 4. There is a green marble in the box if and only if there is a purple marble in the box Conclusion: If there is a blue marble in the box then there is a purple marble in the box.
@@
Statements: 1. There is a purple marble in the box if and only if there is a blue marble in the box. 2. There is a blue marble in the box if and only if there is an olive marble in the box. 3. Either there is an olive marble in the box or else there is a red marble in the box, but not both. 4. Either there is a red marble in the box or else there is a green marble in the box, but not both. Conclusion: If there is a purple marble in the box then there is a green marble in the box.
@@ 2. Statement 2: There is a blue marble in the box if and only if there is an olive marble in the box
Statements: 1. There is a purple marble in the box if and only if there is a blue marble in the box. 2. There is a blue marble in the box if and only if there is an olive marble in the box. 3. Either there is an olive marble in the box or else there is a red marble in the box, but not both. 4. Either there is a red marble in the box or else there is a green marble in the box, but not both. Conclusion: If there is a purple marble in the box then there is a green marble in the box.
@@
Statements: 1. Either there is an olive marble in the box or else there is a brown marble in the box, but not both. 2. There is a brown marble in the box if and only if there is a gray marble in the box. 3. Either there is a gray marble in the box or else there is a maroon marble in the box, but not both. Conclusion: If there is not an olive marble in the box then there is a maroon marble in the box.
@@ 2. Statement 2: There is a brown marble in the box if and only if there is a gray marble in the box
Statements: 1. There is a gray marble in the box if and only if there is a yellow marble in the box. 2. Either there is a yellow marble in the box or else there is an olive marble in the box, but not both. 3. There is an olive marble in the box if and only if there is a black marble in the box. Conclusion: If there is a yellow marble in the box then there is a black marble in the box.
@@
Statements: 1. There is a green marble in the box if and only if there is a purple marble in the box 2. If there is a purple marble in the box then there is a gray marble in the box. 3. Either there is a gray marble in the box or else there is a yellow marble in the box, but not both. Conclusion: Either there is a green marble in the box or else there is a yellow marble in the box, but not both.
@@ 1. First, I’ll start by analyzing the given statements:
Statements: 1. There is a gray marble in the box if and only if there is a yellow marble in the box. 2. Either there is a yellow marble in the box or else there is an olive marble in the box, but not both. 3. There is an olive marble in the box if and only if there is a black marble in the box. Conclusion: If there is a yellow marble in the box then there is a black marble in the box.